OpenStudy (allen27):

Write an equation of a line parallel to y = x - 6 in slope-intercept form that passes through the point (-1, 5). (5 points) y= x - 1 y= x + 5.25 y= -4x + 5 y= -4x - 1.25

OpenStudy (allen27):

@welshfella

OpenStudy (welshfella):

it will have the same slope y = x - 6 recall that the slope of y = mx + b is m so what number multiplies x in y = x - 6?

OpenStudy (welshfella):

- its not there but its 'understood' to be there

OpenStudy (allen27):

i got C last time i tried doing this problem

OpenStudy (welshfella):

wht do you need to multiply x by to give x?

OpenStudy (allen27):

someone just told me to graph it

OpenStudy (welshfella):

gtg right now im sure someone will help..